OpenStudy (brynndelyn):

What is the solution of 16^(-3n+2)=64

OpenStudy (solomonzelman):

Hint: 4^3=64, and hence 16^(3/2)=64

OpenStudy (brynndelyn):

Now I'm even more confused...

OpenStudy (jdoe0001):

hint: \(\bf 16^{-3n+2}=64\qquad \begin{cases} 64=2\cdot 2\cdot 2\cdot 2\cdot 2\cdot 2\to 2^6\\ 16=2\cdot 2\cdot 2\cdot 2\to 2^4 \end{cases} \\ \quad \\ 16^{-3n+2}=64\implies (2^4)^{-3n+2}=(2^6)\)

OpenStudy (brynndelyn):

So you would need to get similar bases then solve? That's so much easier than what I was thinking of

OpenStudy (jdoe0001):

well... yes thus \(\bf 16^{-3n+2}=64\implies (2^4)^{-3n+2}=(2^6) \\ \quad \\ 2^{2(-3n+2)}=2^6\implies 2(-3n+2)=6\)

OpenStudy (brynndelyn):

And that's what I got when I solved, n=6. Thanks!

OpenStudy (jdoe0001):

hmmm

OpenStudy (brynndelyn):

@jdoe0001 Could you help me with another problem?

OpenStudy (jdoe0001):

\(\bf 2^{2(-3n+2)}=2^6\implies 2(-3n+2)=6\implies -6n+4=6\) don't forget to distribute

OpenStudy (jdoe0001):

or you could divide by 2 first either way, is not 6

OpenStudy (jdoe0001):

but you can always post anew for another one, sure more eyes, and if I dunno, someone else may know :)

OpenStudy (brynndelyn):

Are you sure the answer isn't 6?

OpenStudy (jdoe0001):

well.. let's check it let us set n = 6 so we know that 2(-3n+2) =6 so n = 6 thus 2( -3(6) +2 ) =6 (-3(6) + 2) = 3 -18 + 2 = 3 -16 \(\ne\) 3

OpenStudy (anonymous):

Refer to the solution calculation using the Mathematica 9 program.